1) The yield on a 10 year $10,000 par T-bond with 8% semiannual payment coupon and a $10,850 price is:

A. 6.81%

B. 3.41%

C. 8%

D. 7.97%

E.  none of these

2) A 10 year T-bond with semiannual coupon can be stripped into how many separate securities?

Please show work
Answer #1

Semi annual coupon =8%*10000/2 =400
Par Value =10000
Price =10850
Number of Periods =10*2 =20

YTM using Excel formula =2*RATE(20,400,-10850,10000) =6.81% (Option a is correct option)